# Multi-modal Generative AI: Multi-modal LLMs, Diffusions and the Unification

### v2

@@ Abstract
Multi-modal generative AI (Artificial Intelligence) has attracted increasing attention from both academia and industry. Particularly, two dominant families of techniques have emerged: i) Multi-modal large language models (LLMs) demonstrate impressive ability for multi-modal understanding; and ii) Diffusion models exhibit remarkable multi-modal powers in terms of multi-modal generation. Therefore, this paper provides a comprehensive overview of multi-modal generative AI, including multi-modal LLMs, diffusions, and the unification for understanding and generation. To lay a solid foundation for unified models, we first provide a detailed review of both multi-modal LLMs and diffusion models respectively, including their probabilistic modeling procedure, multi-modal architecture design, and advanced applications to image/video LLMs as well as text-to-image/video generation. Furthermore, we explore the emerging efforts toward unified models for understanding and generation. To achieve the unification of understanding and generation, we investigate key designs including autoregressive-based and diffusion-based modeling, as well as dense and Mixture-of-Experts (MoE) architectures. We then introduce several strategies for unified models, analyzing their potential advantages and disadvantages. In addition, we summarize the common datasets widely used for multi-modal generative AI pretraining. Last but not least, we present several challenging future research directions which may contribute to the ongoing advancement of multi-modal generative AI.
Index Terms:
Multi-modal Generative AI, Multi-modal Large Language Model, Diffusion Model, Unified Understanding and Generation
@@ I Introduction
Multi-modal generative AI (Artificial Intelligence) has received increasing attention recently with the advent of (multi-modal) large language models (LLMs) and diffusion models. Two typical models of multi-modal generative AI are GPT-4V [1] and Sora [2] from OpenAI, which have produced great impacts on both academia and industry. To compare GPT-4V and Sora in terms of functionality, GPT-4V targets multi-modal understanding and Sora aims at visual generation — GPT-4V enables the LLM to understand visual input via generating relevant texts, while Sora serves as a text-to-video generation model which outputs visual signals given textual input. To make comparisons in terms of probabilistic modeling, GPT-4V is a multi-modal LLM with autoregressive probabilistic modeling, while Sora is a multi-modal video generation model with diffusion denoising modeling.
@@
To answer these questions, we conduct deep and comprehensive discussions of multi-modal generative AI in this paper, whose overall organization is illustrated in Fig. 1. Specifically, we first present a systematic review of existing works on multi-modal LLM (Sec. II) and multi-modal diffusion (Sec. 4), covering mathematical preliminaries, model architectures, fusion strategies, recent advances, and applications. Then we present our insights on unified models for simultaneous understanding and generation in Sec. IV. Besides, we further summarize video/visual-language datasets for multi-modal generative AI pretraining in Sec. V. Last but not least, we provide future directions that deserve further investigation for multi-modal generative AI.
We would like to point out that although several insightful surveys have been conducted on multi-modal understanding [5, 6, 7], visual generation [8, 9, 10], and both [11, 12], this work differs from them in comprehensive discussions on models for the unification of understanding and generation in addition to reviewing them separately, thus aiming to bridge the gap between understanding and generation in a unified architecture. We highlight recent advances, categorize existing approaches, introduce related datasets, and share insights for future directions. In summary, we make the following contributions.
@@ II Multi-modal LLM for Understanding
Multi-modal LLMs have recently become dominant in the field of visual understanding. In this section, we will review the literature on the multi-modal LLMs.
@@ II-A1 LLM Autoregressive Probabilistic Modeling
The core component of multi-modal LLMs is the LLM, which receives the multi-modal input including the user’s instructions, questions, and visual information, and then outputs the answers to the user in a text-generation form. The LLM is basically an autoregressive model that tries to predict the next word based on all the previous words, as shown in Eq. (1).
@@
where $\theta_{L}$ denotes the parameters of the LLM, which is generally composed of several layers of transformers [13]. Note that LLM can only receive the text tokens as its input, the next important problem for multi-modal LLM is how to enable LLM to understand the visual information. To tackle the problem, most existing works [4, 14, 15] try to align the LLM with the visual encoders from vision-language pretraining tasks, such as CLIP [16]. More recently, there have been some attempts [3] to directly transform the images into discrete visual tokens so that the text and visual tokens can be tackled by the autoregressive LLM together. Next, we will introduce preliminaries about vision-language pretraining and visual tokenizers.
@@ II-A2 Vision-Language Pretraining
CLIP consists of two separate encoders: a visual encoder (typically a Vision Transformer [17] or ResNet [18]) and a text encoder (usually a Transformer). Given a batch of image-text pairs, CLIP is trained with a contrastive loss that encourages the embeddings of matched image-text pairs to be close, while pushing apart the embeddings of mismatched pairs.
The pretrained CLIP model has been widely used in multi-modal LLMs to inject visual understanding into the LLMs. Typically, visual features extracted by the CLIP image encoder are projected into the LLM’s input space through a learned adapter or alignment module [4]. This allows the LLM to reason over both linguistic and visual information in a unified manner.
@@ II-A3 Visual Tokenizer
Inspired by language models where each word is tokenized by a discrete tokenizer, a series of works also transform images into discrete tokens. Typical visual tokenizers include the VQ-VAEs [19, 20] and VQGANs [21, 22], whose overall framework is shown in Fig. 2. We will begin our discussion with VQ-VAE. Basically, VQ-VAE works like an auto-encoder, which has an encoder $E(\cdot)$ and a decoder $D(\cdot)$. Given an image $x$, VQ-VAE first encodes it with an encoder $E(\cdot)$ into a lower-dimensional continuous vector $E(x)$. Then, the continuous vector is discretized using a codebook $Z=\{z_{k}\}_{k=1}^{K}$. The codebook functions similarly to a word embedding table in NLP, where $K$ corresponds to the vocabulary size, and each $z_{k}\in\mathbb{R}^{n_{c}}$ represents a visual prototype analogous to a word embedding. With the encoded vector $E(x)$ and the codebook $Z$, we obtain a discrete representation $z_{q}$ of the image by finding the nearest neighbor of $E(x)$ in $Z$, and use it to reconstruct the image with the decoder: $\hat{x}=D(z_{q})$. This provides a way to convert between images and discrete tokens.
Compared to VQ-VAEs, VQGAN [21, 22] utilizes a GAN perceptual loss to replace the L2 reconstruction loss, which helps to learn a rich codebook. We use a simple example to illustrate the process of tokenization. If we have an input image of size $H\times W\times 3$, after the encoder $E$, we obtain a lower-dimension vector $E(x)$ of size $h\times w\times n_{c}$, where $h<H$ and $w<W$ and $n_{c}$ denote the dimensions of the code. This means we can obtain $h\times w$ vectors of dimension $n_{c}$, and for each vector, we will find its nearest neighbor in the code book for discretization so that we will finally obtain a discrete sequence of length $h\times w$ to represent the image.
Remark. On the one hand, VQGAN and VQ-VAE can be used as visual tokenizers to transform an image into discrete tokens, which enables it to be received by LLMs for visual understanding. On the other hand, they can be used for compressing an image into a lower-dimensional space, which motivates the well-known latent diffusion model (LDM) [23].
@@ II-B Multi-modal LLM Architectures
We categorize existing multi-modal LLM architectures into two branches, the early-fusion architectures and alignment architectures, as shown in Fig. 3. Most existing works [4, 15, 14] adopt the alignment architecture, which aims to align the vision model from the vision-language pretraining with the pretrained LLM. This branch of models relies on the vision-language pretraining to understand the visual input. After obtaining the embedding of the image, an alignment module such as a projector [4], or Q-Former [24] is used to align the image embedding with the LLM space. To train the alignment module, some text-image or text-video pairs are required to input the model. A typical way to align is to make the LLM output the caption of an image given an image embedding. In contrast, as shown on the right of Fig. 3, the early-fusion architectures [3, 25] do not rely on a pretrained vision model to obtain the semantics of the input image. Instead, similar to NLP where each word is mapped to a token, the early-fusion architecture maps each visual input into visual tokens through a visual tokenizer. Then a multi-modal autoregressive language model will receive the mixed text and visual tokens, and output the user’s desired answers.
@@ II-C Image LLM
We will follow the multi-modal LLM architectures section, and elaborate on the latest advancement of image LLM.
@@ II-C1 Alignment-Architecture Image LLM
a) Vision Encoder is a module that extracts crucial information from images. Common generic vision encoders include ResNet [26], the CLIP-ViT encoder [16], and ImageBind [27]. ResNet and CLIP are pretrained on image-text modalities, while ImageBind aligns embeddings from six modalities into a shared space, enabling vision encoders to capture richer information.
b) Alignment Module, also named projector, adapter, etc., aims to mitigate the gap between image features and lexical word tokens and further fuse two modalities. LLaVA [28] adopts a simple but effective linear projection to convert image features into word token embedding space and then it concatenates image tokens and word tokens. Such alignment only involves image transformation, limiting interaction with texts, and is not flexible in the visual token number. Resampler [29] technique maps varying-size features to a fixed number of tokens. BLIP-2 [24] and MiniGPT-4 [30] employ Q-former [24] before linear projections to reduce tokens. Q-former incorporates text semantics and models the interaction between image features and text inputs with learnable queries to enhance the most useful visual content for LLM. Some works focus on preserving locality during projection, such as Honeybee [31], which introduces a locality-enhanced projector to maintain spatial structure. Others prioritize efficiency, like TokenPacker [32], which adopts a coarse-to-fine strategy to compress visual tokens while retaining important details.
@@ II-C2 Early-fusion Architecture Image LLM
Pioneer work Fuyu [33] adopts linear projections on image patches in spatial order and trains a transformer decoder taking the visual and word token sequence as input. Despite limited performance, it reveals a new technical fashion. Google follows this fashion whose Gemini [25] processes the interleaved image and other modalities from the beginning. Chameleon [3] trains an image tokenizer that encodes a 512x512 image into 1024 discrete tokens from a codebook of size 8192. Early-fusion Architecture requires much more computation and it’s more difficult to converge, leaving challenges for future exploration.
@@ II-C3 Challenges in Image LLM
(i) Fine-grained visual concept understanding, where more tokens help encode more detailed information at the cost of causing redundant computation. Chat-UniVi [34] proposes dynamic visual tokens to allocate more computations on important details. An important part of fine-grained understanding is the spatial awareness of object concepts. AnyRef [35] applies RoIAlign to encode regions and designs segment encoder-decoder to learn segmentation from the image LLM’s token outputs, which is similar to OMG-LLaVA [36] who generates pixel- and object-centric visual tokens before projections and decodes segmentation tokens from LLM’s output by OMG-Seg. Different from segmentation supervision, VisionLLM [37] and Virtron [38] use text supervision such as bounding and polygon descriptions by flexible instruction tuning. Fine granularity modeling offers some explanations for LLM. (ii) Hallucination involves errors in objects, attributes, and relations in the forms of judgment or description [39]. Some works [40] try to reduce biases in training data while some mitigate hallucination via improving model characteristics like vision encoders [41] or fusion mechanisms [42]. Human feedbacks [43] also play an important role in reducing hallucination.
@@
Remark. Currently, the alignment architecture still outperforms the early-fusion architecture in multi-modal understanding, e.g., with comparable parameters, the early-fusion architecture Emu3 [44] achieves 75.1 score on VQAv2 [45] benchmark and 58.5 score on MMBench [46] benchmark, while the early-fusion architecture LLAVA-1.6 achieves 86.8 and 67.4 score, respectively. The advantage and disadvantage of the two architectures are as follows: (i) The advantage lies in its capability of utilizing the pretrained knowledge from the vision encoder and LLM. The vision-language pretraining enables the output of the vision encoder to contain semantic meanings. Only the alignment module needs to be trained, which makes this paradigm resource-friendly. (Sometimes other modules are also learnable for better performance.) However, its ability is also limited by the pretrained vision encoder and LLM, e.g., the pretrained CLIP vision encoder often struggles with multiple objects, making the multi-modal LLMs based on CLIP inherit the limitation. (ii) The disadvantage comes from the fact that the early-fusion architecture may have a higher potential, because all its parameters are trained from scratch. However, training from scratch makes the early-fusion architecture face two challenges: (a) a good visual tokenizer needs to be trained and (b) more resources will be needed to train the multi-modal autoregressive model. First, since the visual tokenization process involves compression and discretization, there inevitably exists visual information loss. How to train a tokenizer that contains rich visual information still remains a challenging problem. Second, the visual tokenizers are generally trained with the image reconstruction objective, which in essence belongs to a pixel-level task instead of a semantic-level task. This training strategy requires that the downstream multi-modal LLMs should have an additional ability to learn semantic meanings from the pixel-level information, compared to the original LLMs which are only expected to understand semantic tokens. Therefore, the multi-modal LLMs tend to require much more data for training.

### v3

# Multi-modal Generative AI: Multi-modal LLMs, Diffusions, and the Unification

@@ Abstract
Multi-modal generative AI (Artificial Intelligence) has attracted increasing attention from both academia and industry. Particularly, two dominant families of techniques have emerged: i) Multi-modal large language models (LLMs) demonstrate impressive ability for multi-modal understanding; and ii) Diffusion models exhibit remarkable multi-modal powers in terms of multi-modal generation. Therefore, this paper provides a comprehensive overview of multi-modal generative AI, including multi-modal LLMs, diffusions, and the unification for understanding and generation. To lay a solid foundation for unified models, we first provide a detailed review of both multi-modal LLMs and diffusion models, respectively, including their probabilistic modeling procedure, multi-modal architecture design, and advanced applications to image/video LLMs as well as text-to-image/video generation. Furthermore, we explore the emerging efforts toward unified models for understanding and generation. To achieve the unification of understanding and generation, we investigate key designs including autoregressive-based and diffusion-based modeling, as well as dense and Mixture-of-Experts (MoE) architectures. We then introduce several strategies for unified models, analyzing their potential advantages and disadvantages. In addition, we summarize the common datasets widely used for multi-modal generative AI pretraining. Last but not least, we present several challenging future research directions that may contribute to the ongoing advancement of multi-modal generative AI.
@@ I Introduction
Multi-modal generative AI (Artificial Intelligence) has received increasing attention recently with the advent of (multi-modal) large language models (LLMs) and diffusion models. Two typical models of multi-modal generative AI are GPT-4V [1] and Sora [2] from OpenAI, which have produced great impacts on both academia and industry. To compare GPT-4V and Sora in terms of functionality, GPT-4V targets multi-modal understanding, and Sora aims at visual generation — GPT-4V enables the LLM to understand visual input via generating relevant texts, while Sora serves as a text-to-video generation model which outputs visual signals given textual input. To make comparisons in terms of probabilistic modeling, GPT-4V is a multi-modal LLM with autoregressive probabilistic modeling, while Sora is a multi-modal video generation model with diffusion denoising modeling.
@@
To answer these questions, we conduct deep and comprehensive discussions of multi-modal generative AI in this paper, whose overall organization is illustrated in Fig. 1. Specifically, we first present a systematic review of existing works on multi-modal LLM (Sec. II) and multi-modal diffusion (Sec. 4), covering mathematical preliminaries, model architectures, fusion strategies, recent advances, and applications. Then we present our insights on unified models for simultaneous understanding and generation in Sec. IV. Besides, we further summarize video/visual-language datasets for multi-modal generative AI pretraining in Sec. V. Last, we provide future directions that deserve further investigation for multi-modal generative AI.
In this paper, our scope primarily lies in multi-modal understanding, generation, and their unification. Some concepts widely studied in the field of LLMs, such as in-context learning, post-training techniques (e.g., supervised fine-tuning and reinforcement learning), sparse attention, and positional embeddings, are important but not the main focus of this survey. Readers interested in these topics are referred to related surveys such as [5, 6]. Instead, we focus on recent high-quality works adapted to the multi-modal generative setting, providing a comprehensive overview of the mechanisms that enable multi-modal understanding and generation.
We would like to point out that although several insightful surveys have been conducted on multi-modal understanding [7, 8, 9], visual generation [10, 11, 12, 13, 14], and both [15, 16], this work differs from them in comprehensive discussions on models for the unification of understanding and generation in addition to reviewing them separately, thus contributing to the ongoing advancement of multi-modal generative AI. We highlight recent advances, categorize existing approaches, introduce related datasets, and share insights for future directions. In summary, we make the following contributions.
@@ II Multi-modal LLM for Understanding
Multi-modal LLMs have recently become dominant in the field of understanding. In this section, we will review the literature on the multi-modal LLMs.
@@ II-A1 LLM Autoregressive Probabilistic Modeling
The core component of multi-modal LLMs is the LLM, which receives the multi-modal input, including the user’s instructions, questions, and visual information, and then outputs the answers to the user in a text-generation form. The LLM is basically an autoregressive model that tries to predict the next word based on all the previous words, as shown in Eq. (1).
@@
where $\theta_{L}$ denotes the parameters of the LLM, which is generally composed of several layers of transformers [17]. Note that LLM can only receive the text tokens as its input. The next important problem for multi-modal LLM is how to enable LLM to understand the visual information. To tackle the problem, most existing works [4, 18, 19] try to align the LLM with the visual encoders from vision-language pretraining tasks, such as CLIP [20]. More recently, there have been some attempts [3] to directly transform the images into discrete visual tokens so that the text and visual tokens can be tackled by the autoregressive LLM together. Next, we will introduce preliminaries about vision-language pretraining and visual tokenizers.
@@ II-A2 Vision-Language Pretraining
CLIP consists of two separate encoders: a visual encoder (typically a Vision Transformer [21] or ResNet [22]) and a text encoder (usually a Transformer). Given a batch of image-text pairs, CLIP is trained with a contrastive loss that encourages the embeddings of matched image-text pairs to be close while pushing apart the embeddings of mismatched pairs.
The pretrained CLIP model has been widely used in multi-modal LLMs to inject visual understanding into LLMs. Typically, visual features extracted by the CLIP image encoder are projected into the input space of LLM through a learned adapter or alignment module [4]. This allows LLMs to reason over both linguistic and visual information in a unified manner.
@@ II-A3 Visual Tokenizer
Inspired by language models where each word is tokenized by a discrete tokenizer, a series of works also transform images into discrete tokens. Typical visual tokenizers include the VQ-VAEs [23, 24] and VQGANs [25, 26], whose overall framework is shown in Fig. 2. We will begin our discussion with VQ-VAE. Basically, VQ-VAE works as an auto-encoder with an encoder $E(\cdot)$ and a decoder $D(\cdot)$. Given an image $x$, VQ-VAE first encodes it with an encoder $E(\cdot)$ into a lower-dimensional continuous vector $E(x)$. Then, the continuous vector is discretized using a codebook $Z=\{z_{k}\}_{k=1}^{K}$. The codebook functions similarly to a word embedding table in NLP, where $K$ corresponds to the vocabulary size, and each $z_{k}\in\mathbb{R}^{n_{c}}$ represents a visual prototype analogous to a word embedding. With the encoded vector $E(x)$ and the codebook $Z$, we obtain a discrete representation $z_{q}$ of the image by finding the nearest neighbor of $E(x)$ in $Z$ and use it to reconstruct the image with the decoder: $\hat{x}=D(z_{q})$. This provides a way to convert between images and discrete tokens.
Compared to VQ-VAEs, VQGAN [25, 26] utilizes a GAN perceptual loss to replace the L2 reconstruction loss, which helps to learn a rich codebook. We use a simple example to illustrate the tokenization process. If we have an input image of size $H\times W\times 3$, after the encoder $E$, we obtain a lower-dimension vector $E(x)$ of size $h\times w\times n_{c}$, where $h<H$, $w<W$, and $n_{c}$ denote the dimensions of the code. This means that we can obtain $h\times w$ vectors of dimension $n_{c}$, and for each vector we will find its nearest neighbor in the code book for discretization so that we will finally obtain a discrete sequence of length $h\times w$ to represent the image.
Remark. On the one hand, VQGAN and VQ-VAE can be used as visual tokenizers to transform an image into discrete tokens, which enables it to be received by LLMs for visual understanding. On the other hand, they can be used to compress an image into a lower-dimensional space, which motivates the well-known latent diffusion model (LDM) [27].
@@ II-B Multi-modal LLM Architectures
We categorize existing multi-modal LLM architectures into two branches, the alignment architectures and the early-fusion architectures, as shown in Fig. 3. Most existing works [4, 19, 18] adopt the alignment architecture, which aims to align the vision model from the vision-language pretraining with the pretrained LLM. This branch of models relies on the vision-language pretraining to understand the visual input. After obtaining the embedding of the image, an alignment module such as a projector [4] or Q-Former [28] is used to align the image embedding with the LLM space. To train the alignment module, some text-image or text-video pairs are required to input the model. A typical way to align is to make the LLM output the caption of an image given an image embedding. In contrast, as shown on the right of Fig. 3, the early-fusion architectures [3, 29] do not rely on a pretrained vision model to obtain the semantics of the input image. Instead, similar to NLP, where each word is mapped to a token, the early-fusion architecture maps each visual input into visual tokens through a visual tokenizer. Then, a multi-modal autoregressive language model will receive the mixed text and visual tokens and output the user’s desired answers.
@@ II-C Image LLM
We will follow the multi-modal LLM architectures section and elaborate on the latest advancement of image LLM.
@@ II-C1 Alignment-Architecture Image LLM
a) Vision Encoder is a module that extracts crucial information from images. Common generic vision encoders include ResNet [30], the CLIP-ViT encoder [20], and ImageBind [31]. ResNet and CLIP are pretrained on image-text modalities, while ImageBind aligns embeddings from six modalities into one shared space, enabling vision encoders to capture richer information.
b) Alignment Module, also named projector, adapter, etc., aims to mitigate the gap between image features and lexical word tokens and further fuse two modalities. LLaVA [4] adopts a simple but effective linear projection to convert image features into word token embedding space and then it concatenates image tokens and word tokens. Such alignment only involves image transformation, limiting interaction with texts, and is not flexible in the visual token number. Resampler [32] technique maps varying-size features to a fixed number of tokens. BLIP-2 [28] and MiniGPT-4 [33] employ Q-former [28] before linear projections to reduce tokens. Q-former incorporates text semantics and models the interaction between image features and text inputs with learnable queries to enhance the most useful visual content for LLM. Some works focus on preserving locality during projection, such as Honeybee [34], which introduces a locality-enhanced projector to maintain spatial structure. Others prioritize efficiency, such as TokenPacker [35], which adopts a coarse-to-fine strategy to compress visual tokens while retaining important details.
@@ II-C2 Early-fusion Architecture Image LLM
Pioneer work Fuyu [36] adopts linear projections on image patches in spatial order and trains a transformer decoder taking the visual and word token sequence as input. Despite limited performance, it reveals a new technical fashion. Google follows this fashion, whose Gemini [29] processes the interleaved image and other modalities from the beginning. Chameleon [3] trains an image tokenizer that encodes a 512x512 image into 1024 discrete tokens from a codebook of size 8192. Early-fusion Architecture requires more computation and is more difficult to converge, leaving challenges for future exploration.
@@ II-C3 Challenges in Image LLM
(i) Fine-grained visual concept understanding, where more tokens help encode more detailed information at the cost of causing redundant computation. Chat-UniVi [37] proposes dynamic visual tokens to allocate more computations on important details. An important part of fine-grained understanding is the spatial awareness of object concepts. AnyRef [38] applies RoIAlign to encode regions and designs a segment encoder-decoder to learn segmentation from the image LLM’s token outputs, which is similar to OMG-LLaVA [39], who generates pixel- and object-centric visual tokens before projections and decodes segmentation tokens from LLM’s output by OMG-Seg. Different from segmentation supervision, VisionLLM [40] and Virtron [41] use text supervision such as bounding and polygon descriptions by flexible instruction tuning. Fine granularity modeling offers some explanations for LLM. (ii) Hallucination involves errors in objects, attributes, and relations in the forms of judgment or description [42]. Some works [43] try to reduce biases in training data, while some mitigate hallucination by improving model characteristics such as vision encoders [44] or fusion mechanisms [45]. Human feedbacks [46] also play an important role in reducing hallucination.
@@
Remark. Currently, the alignment architecture still outperforms the early-fusion architecture in multi-modal understanding, e.g., with comparable parameters, the early-fusion architecture Emu3 [47] achieves 75.1 score on VQAv2 [48] benchmark and 58.5 score on MMBench [49] benchmark, while the early-fusion architecture LLAVA-1.6 achieves 86.8 and 67.4 score, respectively. The advantages and disadvantages of the two architectures are as follows: (i) The advantage lies in the capability of utilizing the pretrained knowledge from the vision encoder and LLM. The vision-language pretraining enables the output of the vision encoder to contain semantic meanings. Only the alignment module needs to be trained, which makes this paradigm resource-friendly. (Sometimes other modules are also learnable for better performance.) However, its ability is also limited by the pretrained vision encoder and LLM, e.g., the pretrained CLIP vision encoder often struggles with multiple objects, making the multi-modal LLMs based on CLIP inherit the limitation. (ii) The disadvantage comes from the fact that the early-fusion architecture may have a higher potential, because all its parameters are trained from scratch. However, training from scratch makes the early-fusion architecture face two challenges: (a) a good visual tokenizer needs to be trained, and (b) more resources will be needed to train the multi-modal autoregressive model. First, since the visual tokenization process involves compression and discretization, there inevitably exists visual information loss. How to train a tokenizer that contains rich visual information still remains a challenging problem. Second, the visual tokenizers are generally trained with the image reconstruction objective, which in essence belongs to a pixel-level task instead of a semantic-level task. This training strategy requires that the downstream multi-modal LLMs should have an additional ability to learn semantic meanings from the pixel-level information, compared to the original LLMs, which are only expected to understand semantic tokens. Therefore, multi-modal LLMs tend to require more data for training.